OpenStudy (anonymous):

Knee socks are 2 pair for $2.18, and anklets are 3 pair for $3.15. What of the following is true? A. Knee socks are 9 cence more per pair thank ankelts. B. Knee socks are 9 cence less pair than anklets. C. Knee socks are $1.09 per pair or D. Anklets are $1.09 per pair

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

1.09 +1.09=2.18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Do you need help with fractions

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but I need help with something else

OpenStudy (anonymous):

give me your other questions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The distance between springfield and Summerville is 24 miles. On Marys map this distance is 3 inches. Winterhaven is 16 miles from springfield. On the map, what is the distance between springfield and winterhaven?

OpenStudy (anonymous):

i'm not sure on this one sorry

OpenStudy (anonymous):

there is no answers

OpenStudy (anonymous):

I will try another

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i mean there is no A.,B.,C.,D.

OpenStudy (anonymous):

4x-8 =12

OpenStudy (anonymous):

yah there is a b c and d

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

BRB

OpenStudy (anonymous):

A. 0.8 inches B. 1.3 inches C. 2.0 inches D. 2.4 inches

OpenStudy (anonymous):

I can only do Decimals and fractions like turning fractions into decimals and stuff like that and more

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

sorry!!!

OpenStudy (anonymous):

Give me a different one i will see if i can answer

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x-8=12

OpenStudy (anonymous):

Which of the following is true for the same value as x

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

what is a equation for it

OpenStudy (anonymous):

for how to solve it

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

WHAT IS AN EUQATION FOR THE SAME VALUE FOR X

OpenStudy (anonymous):

sorry i was in the bathroom

OpenStudy (anonymous):

lol

OpenStudy (anonymous):